Describe stearic acid. What is the partial charge characteristic of each carbon atom in the molecule of this acid?
1
Expert's answer
2020-11-26T10:54:10-0500

Stearic acid is a saturated and long-chain fatty acid which entails a 18-carbon backbone"C17H35C000H"

"Let the charge of Carbon be X"

"x+17+H35+ x+ O-2+O-2+H1=0"

"x+17+35+x+-2+-2+1=0"

"2x+17+35+-4+1=0"

"2x +52-4+1=0"

"2x+49=0"

"x=-49\/2"

x=-25

"Thereforef, each Carbon aton will have a charge of -25"